OpenStudy (anonymous):

a body released from the top of a tower of height h takes t second to reach the ground at time 1/2t the height of the body from the ground is?

OpenStudy (anonymous):

4.9 unless i using wrong formula D=.5 * accelaration * time squared .5*9.8

OpenStudy (anonymous):

the ans=3h/4

OpenStudy (anonymous):

4.9 meters All objects fall at same 9.8 m/s^2

OpenStudy (anonymous):

look like typical question

OpenStudy (stormfire1):

3h/4 (75% of the initial height) is right

OpenStudy (anonymous):

first find t in terms of h. For the stone to drop from rest until it reaches the ground : we use \[s= ut + 0.5at^2\] where s=h, u =0, a assumed to be 10 then you will get \[t=\sqrt{h/5}\] After 1/2 t , let the stone travels a distance of x unit still use the formula \[s= ut + 0.5at^2\] \[x= 0 + 0.5(10)(0.5t)^2\] \[x= 5/4 t^2\] sub the value of t you found just now and you will get x=h/4 so it's height from the ground is = h-h/4 = 3h/4

OpenStudy (anonymous):

should be \[x=(5/4)t^2\]

OpenStudy (stormfire1):

I figured :)

OpenStudy (anonymous):

how 5/4???

OpenStudy (anonymous):

not getting that...

OpenStudy (anonymous):

\[x= (0.5)(10)(0.5^2)t^2 \] use your calculator to get the value 5/4

OpenStudy (anonymous):

i didnt see 0.5^2

OpenStudy (anonymous):

thanzzz